Date Ending Inventory (End-of-Year Prices) Price Index December 31, 2017 $ 73,200 100 December 31, 2018...

Date

Ending Inventory
(End-of-Year Prices)

Price
Index

December 31, 2017

$ 73,200 100

December 31, 2018

128,876 116

December 31, 2019

124,608 132

December 31, 2020

140,426 143

December 31, 2021

168,476 154

December 31, 2022

199,520 160


Compute the ending inventory for Nash Company for 2017 through 2022 using the dollar-value LIFO method.

Ending Inventory

2017

2018

2019

$enter a dollar amount

2020

2021

2022
Answer #1

Calculate ending inventory

Ending inventory
2017 73200
2018 128876/1.16 = 111100 73200+37900*1.16 = 117164
2019 124608/1.32 = 94400 73200+21200*1.16 = 97792
2020 140426/1.43 = 98200 97792+3800*1.43 = 103226
2021 168476/1.54 = 109400 103226+11200*1.54 = 120474
2022 199520/1.60 = 124700 120474+15300*1.60 = 144954